a sum of $3600 together with the interest earned is left for 3 years in a bank that pays an interest rate of 6% compounded annually. How much interest is earned (to the nearest cent)?
Answer by stanbon(48557)

a sum of $3600 together with the interest earned is left for 3 years in a bank that pays an interest rate of 6% compounded annually. How much interest is earned (to the nearest cent)?
-------------------------------
A(3) = 3600**(1.06)^3 = $4287.6576
Inerest = 4287.6576-3600 = $687.66
=================================
Cheers,
Stan H.